# Generative Ghosts: Anticipating Benefits and Risks of AI Afterlives

### v2

@@ Abstract
As AI systems quickly improve in both breadth and depth of performance, they lend themselves to creating increasingly powerful and realistic agents, including the possibility of agents modeled on specific people. We anticipate that within our lifetimes it may become common practice for people to create a custom AI agent to interact with loved ones and/or the broader world after death. We call these generative ghosts, since such agents will be capable of generating novel content rather than merely parroting content produced by their creator while living. In this paper, we first discuss the design space of potential implementations of generative ghosts. We then discuss the practical and ethical implications of generative ghosts, including potential positive and negative impacts on individuals and society. Based on these considerations, we lay out a research agenda for the AI and HCI research communities to empower people to create and interact with AI afterlives in a safe and beneficial manner.
Introduction
@@
The past few years have brought incredible growth in the capabilities of generative AI models, particularly large language models (LLMs) such as GPT-4 (OpenAI 2023a), Palm 2 (Anil et al. 2023), and Llama 2 (Touvron et al. 2023), though there has also been incredible progress in generative AI for the production of images (Ramesh et al. 2022), video (Singer et al. 2022), and audio (Borsos et al. 2023), as well as a new generation of multimodal models (Yang et al. 2023; Google DeepMind 2023) that combine functionality across several media categories. These models, in turn, have given rise to new types of generative agents (Park et al. 2023), simulacra that can produce believable human behaviors, including capabilities such as memory and planning. While still in their infancy, generative agents and related technologies are likely to increase in fidelity and popularity as underlying model capabilities improve and compute costs drop. For instance, in November 2023 OpenAI released GPTs (OpenAI 2023b), a no-code interface for people to develop agentic AIs.
@@
As AI models increase the set of human capabilities they can faithfully reproduce (Morris et al. 2023; Bubeck et al. 2023), societal change is inevitable. For instance, experts anticipate that powerful AI systems may profoundly change disparate areas of society such as the labor market (Eloundou et al. 2023), the education system (Kasneci et al. 2023), the pursuit of scientific knowledge (Morris 2023), and criminal activities (Ferrara 2023). In this paper, we discuss how advances in AI might change personal and cultural practices around death and dying.
We introduce the concept of generative ghosts, AI agents that represent a deceased person, and discuss why we anticipate such representations will become popular within our lifetimes. We explore the design space of possible instantiations of generative ghosts and consider both the benefits that might lead to their adoption and the practical and ethical concerns such technology may introduce.
@@
Our contributions include: (1) identifying and characterizing an emerging phenomenon of creating “generative ghosts” to represent the deceased, (2) introducing a taxonomy of design dimensions and analysis of potential benefits and harms that can be used to support future empirical research and motivate fieldwork. By characterizing this emerging trend, highlighting potential risks to be averted, and creating a framework for future investigation, we aim to ensure that future technical and sociotechnical systems will maximize the potential benefits of “AI afterlives” while minimizing potential risks.
Related Work
@@
Post-Mortem Technology
@@
Throughout history, people have turned to technology to remember, memorialize, and even interact with the dead. Gravestones and other burial markers can be traced nearly back to 3000 B.C.E. (Taylor 2001). Obituaries in the U.S., while dating back to the 16th century, became more common during the 19th century in part due to the U.S. Civil War (Hume 2000) – an event that also brought embalming into favor. Even the mediums of the Spiritualism movement in the late 19th and early 20th century turned to telegraphs, radio-wave detectors, and later wireless radio in their attempts to detect the presence of and communicate with the dead (National Science and Media Museum 2022).
@@
During the earliest days of the World Wide Web, when people would create personal Home Pages describing their lives and family, it was routine for people to dedicate a page to the memory of a family member, often a deceased parent or family pet (Brubaker 2015). Online graveyards, websites specifically dedicated to the memorialization of the dead, soon followed (Roberts and Vidal 2000; Roberts 2004). The Virtual Memorial Garden is the earliest documented example (Roberts and Vidal 2000). Created in 1996, it featured a collection of brief obituaries authored by loved ones, capturing their grief, the circumstances around the death, and loved ones’ guilt.
@@
Even as people adopted digital technology to memorialize the dead, scholars in human-computer interaction and social computing began to note how mortality is overlooked by technology design and have argued for increased attention to end-of-life and mortality (Massimi and Charise 2009; Massimi et al. 2010; Odom et al. 2010; Bell 2006), often termed “thanatosensitive design” (Massimi and Charise 2009). Scholars have studied the intersection of technology and mortality across diverse contexts, including digital heirlooms (e.g., (Odom et al. 2012; Beuthel and Fuchsberger 2022)), communal rituals (e.g., (Häkkilä, Colley, and Kalving 2019; Uriu et al. 2021)), online memorials (e.g., (Gulotta et al. 2014; Moncur and Kirk 2014; Massimi 2013)), digital legacy (e.g., (Gulotta et al. 2016, 2014; Gulotta, Kelliher, and Forlizzi 2017; Doyle and Brubaker 2023)), and family archives (e.g., (Kaye et al. 2006)). Yet, as Doyle and Brubaker summarize (Doyle and Brubaker 2024), end-of-life scenarios also present new privacy challenges (e.g., (Holt, Nicholson, and Smeddinck 2021; Locasto, Massimi, and DePasquale 2011)), challenges that result from shifting motivations at different life stages (Chen, Vitale, and McGrenere 2021; Thomas and Briggs 2014), and challenges due to differences between user expectations and platform functionality (Gach and Brubaker 2020).
@@
In HCI, extensive attention has been given to digital legacies, the collection of materials that carry values and meaning that are passed down and/or otherwise continue to represent the deceased after their death. A recent literature review of the field by Doyle and Brubaker (Doyle and Brubaker 2023) identified four foci of this scholarship: digital identity (i.e., how legacies can continue to represent the deceased in intentional and unintentional ways), engagement with digital legacies (i.e., studies on use and user perceptions), putting to rest (concerned with issues around preservation, disposition, and disposal), and the integration of technology into existing legacy practices (Doyle and Brubaker 2023).
@@
In most cases, it is the digital traces people leave behind that give rise to legacy crafting and memorialization opportunities. The sources of legacy content vary – from burner accounts (Gulotta et al. 2016) to personal archives (She et al. 2021; Kaye et al. 2006) – but social media content (especially memorialized profiles) has received the most scholarly attention (e.g., Brubaker et al. 2014; Gach and Brubaker 2020, 2021; Brubaker and Callison-Burch 2016; Brubaker and Hayes 2011; Brubaker, Hayes, and Mazmanian 2019; Getty et al. 2011; Mori et al. 2012; Sofka 2009, 2017; Arnold et al. 2017; Nansen et al. 2019).
@@
As Doyle and Brubaker (2023) summarize, the persistence of identities has benefits and challenges: Digital legacies can facilitate healthy grieving (She et al. 2021; c.f., Brubaker, Hayes, and Dourish 2013; Brubaker et al. 2012) and maintain connections to the deceased (Brubaker and Hayes 2011; Getty et al. 2011; Gulotta et al. 2014). However, receiving what is often a large and uncurated set of content can be overwhelming for loved ones (Holt, Nicholson, and Smeddinck 2021), and may provide (for better or worse) an uncensored version of the deceased (Gulotta et al. 2013). Loved ones may also find themselves overwhelmed as they become the contact point for the deceased’s various online networks (Brubaker et al. 2014; Brubaker, Hayes, and Dourish 2013).
@@
Across many studies, participants describe uncertainty around the use, care, and understanding of digital legacies. Ambiguities often arise when data created for one purpose – e.g., personal archives (Kaye et al. 2006; Holt, Nicholson, and Smeddinck 2021) or interpersonal communication (Moncur and Kirk 2014; Thomas and Briggs 2014; Gach and Brubaker 2021, 2020; Gulotta et al. 2014) – are repurposed post-mortem. In contrast to traditional legacies where there are familiar norms and practices, Pfister (2017) argues that common norms around digital legacy have not yet emerged, an issue that AI will certainly exacerbate.
AI Afterlives
@@
In recent years, a handful of tech-savvy individuals have attempted to posthumously create interactive memorializations of loved ones by ingesting their digital (or digitized) content such as emails, journals, videos, photographs, or other autobiographical media. In her graphic novel Artificial: A Love Story, Amy Kurzweil 2023 documents how her father, futurist Ray Kurzweil, created a chatbot to embody the memory of his deceased father, Fred Kurzweil, in a form he dubs “Fredbot”(Veltman 2023). Fredbot interactively responds to questions from his descendants, but only by sharing exact quotes from the materials Fred left behind such as letters digitized by his family. In another well-publicized incident, the engineer Eugenia Kuyda created an app to preserve the memory of her best friend, Roman, who died unexpectedly in an accident, training a neural network on text messages her friend had sent her to create a bot named, suitably, “Roman.” Unlike Fredbot, which was available only to immediate family, the Roman bot was made available on social media and app stores for public interaction, resulting in mixed reactions from friends and family of the deceased (Newton 2016).While Fredbot and Roman bot represented private citizens, AI has also been used to re-create public figures in various ways. For instance, in November 2024, the surviving members of The Beatles released a new song, “Now and Then,” (Beatles 2023) using AI to enable the deceased John Lennon to sing along with his living bandmates (Savage 2023). The musician Laurie Anderson collaborated with the University of Adelaide’s Australian Institute for Machine Learning to create a chatbot based on her longtime partner, deceased musician Lou Reed, which, in addition to conversing, generates new music lyrics in Reed’s style (Andereson 2024). In early 2024, gun-control activists in the used AI to re-create the voices of victims of gun violence, to create a visceral and poignant plea to U.S. legislators (Luscombe 2024).
@@
Meanwhile, there are a handful of start-up companies that aim to give people the ability to create their own AI afterlives while they are still alive. Re;memory (rememory.deepbrain.io/en) is an offering from DeepBrain AI that professes to create an interactive virtual representation after a seven-hour filming and interview session. They specifically advertise this service as a way that one can proactively create a rich memory that friends and family can engage with after one dies. HereAfter (https://www.hereafter.ai) provides an app that interviews a user with the goal of proactively creating a posthumous digital representation. Friends and relatives can interact with a chatbot-based representation of their loved one who can provide photos and voice recordings of memories and life events. Companies offering ”digital immortality” are increasingly popular in China (Yang 2024).
@@
In contrast with Re;memory and HereAfter, which aim to empower users to preserve their memories before their deaths, some start-ups purport to use AI to provide an experience more akin to resurrection or reanimation (i.e., without the planning or consent of the deceased party). For instance, Character.AI (https://beta.character.ai/) uses LLMs to enable a chat experience with deceased public figures for entertainment or educational purposes (e.g., chatting with a bot representing William Shakespeare), and the Musée D’Orsay in Paris recently contracted with a start-up company to develop a chatbot meant to represent the artist Vincent van Gogh by ingesting his writings (Small 2023). Beyond public figures, Project December (projectdecember.net/) is a venture that allows end-users to upload content from a deceased loved one (Fagone 2021) in order to “simulate the dead” via a customized chatbot.
@@
While the examples shared above demonstrate the diversity of emerging approaches, a key trend in post-mortem AI is around griefbots. Griefbots (Jiménez-Alonso and de Luna 2022; van der Vorst and Kamp 2022) or deathbots (Lindemann 2022) are typically described as chatbots (predating modern large language model technologies) designed to allow the bereaved to “converse” with the deceased, typically created posthumously by a third party who leverages general-purpose text the deceased created during their life. The aforementioned “Fredbot” and “Roman” would be considered examples of griefbots. Because of their third-party origins, griefbots raise controversial ethical issues regarding privacy and consent (Grandinetti, DeAtley, and Bruinsma 2020).
@@
Generative ghosts can be considered an extension of the concept of a griefbot, but not all griefbots would be considered generative ghosts. Generative ghosts go beyond the traditional definition of griefbots in several ways: they always include the ability to generate novel content in-character (whereas many griefbots only regurgitate content from a training corpus), potentially evolve over time, and possess agentic (Park et al. 2023; Shavit et al. 2023) capabilities such as the ability to participate in the economy or perform other complex tasks with limited oversight. Notably, our framework for generative ghosts envisions the possibility of first-party (rather than only third-party) creation as part of an end-of-life planning process. Also, generative ghosts may be designed to represent non-human entities (such as pets), and in many cases might exist pre-mortem as a generative clone. Finally, in contrast to griefbots, generative ghosts may be created for reasons other than to support the bereaved in their grieving process. The following section introduces more details about the design space of generative ghosts.
Generative Ghosts: A Design Space

### v3

@@ Abstract.
As AI systems quickly improve in both breadth and depth of performance, they lend themselves to creating increasingly powerful and realistic agents, including the possibility of agents modeled on specific people. We anticipate that within our lifetimes it may become common practice for people to create a custom AI agent to interact with loved ones and/or the broader world after death; indeed, the past year has seen a boom in startups purporting to offer such services. We call these generative ghosts, since such agents will be capable of generating novel content rather than merely parroting content produced by their creator while living. In this paper, we reflect on the history of technologies for AI afterlives, including current early attempts by individual enthusiasts and by startup companies to create generative ghosts. We then introduce a novel design space detailing potential implementations of generative ghosts, and use this taxonomy to ground discussion of the practical and ethical implications of various approaches to designing generative ghosts, including potential positive and negative impacts on individuals and society. Based on these considerations, we lay out a research agenda for the AI and HCI research communities to better understand the risk/benefit landscape of this novel technology so as to ultimately empower people who wish to create and interact with AI afterlives to do so in a safe and beneficial manner.
AI, AI agents, Generative AI, HCI, digital afterlife, digital legacy, post-mortem AI, post-mortem data management, end-of-life planning, death, griefbots
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
@@
The past few years have brought incredible growth in the capabilities of generative AI models, particularly large language models (LLMs) such as GPT-4 (OpenAI, 2023a), Palm 2 (Anil et al., 2023), and Llama 2 (Touvron et al., 2023), though there has also been incredible progress in generative AI for the production of images (Ramesh et al., 2022), video (Singer et al., 2022), and audio (Borsos et al., 2023), as well as a new generation of multimodal models (Yang et al., 2023; Google DeepMind, 2023) that combine functionality across several media categories. These models, in turn, have given rise to new types of generative agents (Park et al., 2023), simulacra that can produce believable human behaviors, including capabilities such as memory and planning. While still in their infancy, generative agents and related technologies are likely to increase in fidelity and popularity as underlying model capabilities improve and compute costs drop. For instance, in November 2023 OpenAI released GPTs (OpenAI, 2023b), a no-code interface for people to develop agentic AIs.
@@
As AI models increase the set of human capabilities they can faithfully reproduce (Morris et al., 2023; Bubeck et al., 2023), societal change is inevitable. For instance, experts anticipate that powerful AI systems may profoundly change disparate areas of society such as the labor market (Eloundou et al., 2023), the education system (Kasneci et al., 2023), the pursuit of scientific knowledge (Morris, 2023), and criminal activities (Ferrara, 2023). In this paper, we discuss how advances in AI might change personal and cultural practices around death and dying.
We introduce the concept of generative ghosts, AI agents that represent a deceased person, and discuss why we anticipate such representations may become increasingly popular. We share examples from recent media reports of individuals and startups experimenting with attempting to create generative ghosts, a trend that seems to be accelerating, particularly in East Asia. We introduce a design space of possible instantiations of generative ghosts, and use this taxonomy to structure an analysis of how design choices might lead to practical and ethical concerns as well as potentially beneficial outcomes.
@@
Our contributions include: (1) identifying and characterizing an emerging phenomenon of creating “generative ghosts” to represent the deceased, (2) introducing a taxonomy of design dimensions and analysis of potential benefits and harms that can be used to support future empirical research and motivate fieldwork. By characterizing this emerging trend, highlighting potential risks to be averted, and creating a framework for future investigation, we aim to ensure that any future technical and sociotechnical systems will maximize the potential benefits of “AI afterlives” while minimizing potential risks.
2. Related Work
@@
2.1. Post-Mortem Technology
@@
Throughout history, people have turned to technology to remember, memorialize, and even interact with the dead. Gravestones and other burial markers can be traced nearly back to 3000 B.C.E. (Taylor, 2001). Obituaries in the U.S., while dating back to the 16th century, became more common during the 19th century in part due to the U.S. Civil War (Hume, 2000) – an event that also brought embalming into favor. Even the mediums of the Spiritualism movement in the late 19th and early 20th century turned to telegraphs, radio-wave detectors, and later wireless radio in their attempts to detect the presence of and communicate with the dead (National Science and Media Museum, 2022).
@@
During the earliest days of the World Wide Web, when people would create personal Home Pages describing their lives and family, it was routine for people to dedicate a page to the memory of a family member, often a deceased parent or family pet (Brubaker, 2015). Online graveyards, websites specifically dedicated to the memorialization of the dead, soon followed (Roberts and Vidal, 2000; Roberts, 2004). The Virtual Memorial Garden is the earliest documented example (Roberts and Vidal, 2000). Created in 1996, it featured a collection of brief obituaries authored by loved ones, capturing their grief, the circumstances around the death, and loved ones’ guilt.
@@
Even as people adopted digital technology to memorialize the dead, scholars in human-computer interaction and social computing began to note how mortality is overlooked by technology design and have argued for increased attention to end-of-life and mortality (Massimi and Charise, 2009; Massimi et al., 2010; Odom et al., 2010; Bell, 2006), often termed “thanatosensitive design” (Massimi and Charise, 2009). Scholars have studied the intersection of technology and mortality across diverse contexts, including digital heirlooms (e.g., (Odom et al., 2012; Beuthel and Fuchsberger, 2022)), communal rituals (e.g., (Häkkilä et al., 2019; Uriu et al., 2021)), online memorials (e.g., (Gulotta et al., 2014; Moncur and Kirk, 2014; Massimi, 2013)), digital legacy (e.g., (Gulotta et al., 2016, 2014, 2017; Doyle and Brubaker, 2023)), and family archives (e.g., (Kaye et al., 2006)). Yet, as Doyle and Brubaker summarize (Doyle and Brubaker, 2024), end-of-life scenarios also present new privacy challenges (e.g., (Holt et al., 2021; Locasto et al., 2011)), challenges that result from shifting motivations at different life stages (Chen et al., 2021; Thomas and Briggs, 2014), and challenges due to differences between user expectations and platform functionality (Gach and Brubaker, 2020).
@@
In HCI, extensive attention has been given to digital legacies, the collection of materials that carry values and meaning that are passed down and/or otherwise continue to represent the deceased after their death. A recent literature review of the field by Doyle and Brubaker (Doyle and Brubaker, 2023) identified four foci of this scholarship: digital identity (i.e., how legacies can continue to represent the deceased in intentional and unintentional ways), engagement with digital legacies (i.e., studies on use and user perceptions), putting to rest (concerned with issues around preservation, disposition, and disposal), and the integration of technology into existing legacy practices (Doyle and Brubaker, 2023).
@@
In most cases, it is the digital traces people leave behind that give rise to legacy crafting and memorialization opportunities. The sources of legacy content vary – from burner accounts (Gulotta et al., 2016) to personal archives (She et al., 2021; Kaye et al., 2006) – but social media content (especially memorialized profiles) has received the most scholarly attention (e.g., (Brubaker et al., 2014; Gach and Brubaker, 2020, 2021; Brubaker and Callison-Burch, 2016; Brubaker and Hayes, 2011; Brubaker et al., 2019; Getty et al., 2011; Mori et al., 2012; Sofka, 2009, 2017; Arnold et al., 2017; Nansen et al., 2019)).
@@
The persistence of identities has benefits and challenges: Digital legacies can facilitate healthy grieving (She et al., 2021; Brubaker et al., 2013, 2012) and maintain connections to the deceased (Brubaker and Hayes, 2011; Getty et al., 2011; Gulotta et al., 2014). However, receiving a large and uncurated set of content can be overwhelming for loved ones (Holt et al., 2021), and may provide (for better or worse) an uncensored version of the deceased (Gulotta et al., 2013). Loved ones may also find themselves overwhelmed as they become the contact point for the deceased’s various online networks (Brubaker et al., 2014, 2013).
@@
Across many studies, participants describe uncertainty around the use, care, and understanding of digital legacies. Ambiguities often arise when data created for one purpose – e.g., personal archives (Kaye et al., 2006; Holt et al., 2021) or interpersonal communication (Moncur and Kirk, 2014; Thomas and Briggs, 2014; Gach and Brubaker, 2021, 2020; Gulotta et al., 2014) – are repurposed post-mortem. In contrast to traditional legacies where there are familiar norms and practices, Pfister (Pfister, 2017) argues that common norms around digital legacy have not yet emerged, an issue that AI will certainly exacerbate.
2.2. AI Afterlives
@@
A handful of tech-savvy individuals have attempted to create interactive memorializations of loved ones by ingesting their digital (or digitized) content such as emails, journals, videos, photographs, or other autobiographical media. In her graphic novel Artificial: A Love Story, Amy Kurzweil (Kurzweil, 2023) documents how her father, futurist Ray Kurzweil, created a chatbot to embody the memory of his deceased father, Fred Kurzweil, in a form he dubs “Fredbot” (Veltman, 2023). Fredbot interactively responds to questions from his descendants, but only by sharing exact quotes from the materials Fred left behind such as letters digitized by his family. In another well-publicized incident, the engineer Eugenia Kuyda created an app to preserve the memory of her best friend, Roman, who died unexpectedly in an accident, training a neural network on text messages her friend had sent her to create a bot named, suitably, “Roman.” Unlike Fredbot, which was available only to immediate family, the Roman bot was made available on social media and app stores for public interaction, resulting in mixed reactions from friends and family of the deceased (Newton, 2016). While Fredbot and Roman bot represented private citizens, AI has also been used to re-create public figures in various ways. For instance, in November 2023, the surviving members of The Beatles released a new song, “Now and Then,” (Beatles, 2023) using AI to enable the deceased John Lennon to sing along with his living bandmates (Savage, 2023). The musician Laurie Anderson collaborated with the University of Adelaide’s Australian Institute for Machine Learning to create a chatbot based on her longtime partner, deceased musician Lou Reed, which, in addition to conversing, generates new music lyrics in Reed’s style (Andereson, 2024). In early 2024, gun-control activists in the used AI to re-create the voices of victims of gun violence, to create a visceral and poignant plea to U.S. legislators (Luscombe, 2024).
@@
Meanwhile, there are a handful of start-up companies that aim to give people the ability to create their own AI afterlives while they are still alive. Re;memory (rememory.deepbrain.io/en) is an offering from DeepBrain AI that professes to create an interactive virtual representation after a seven-hour filming and interview session. They specifically advertise this service as a way that one can proactively create a rich memory that friends and family can engage with after one dies. HereAfter (https://www.hereafter.ai) provides an app that interviews a user with the goal of proactively creating a posthumous digital representation. Friends and relatives can interact with a chatbot-based representation of their loved one who can provide photos and voice recordings of memories and life events. Early attempts at generative ghost technologies seem to already be entering the mainstream in East Asia, where ”communicating” (via non-technological means) with deceased ancestors is already a cultural norm for many (Yang, 2024a). Companies offering ”digital immortality” are increasingly popular in China, serving thousands of customers (Yang, 2024b; Feng, 2024; Yang, 2024a). In South Korea, 19 million people have streamed an emotional video of a bereaved mother, Jang Ji-Sun, interacting with a virtual reality representation of her deceased young daughter, Na-yeon, custom-created for her by a media company (Chesler, 2022).
@@
In contrast with Re;memory and HereAfter, which aim to empower users to preserve their memories before their deaths, some start-ups purport to use AI to provide an experience more akin to resurrection or reanimation (i.e., without the consent of the deceased party). For instance, Character.AI (https://beta.character.ai/) uses LLMs to simulate chatting with deceased public figures for entertainment or education (e.g., a bot representing William Shakespeare), and the Musée D’Orsay in Paris recently contracted with a start-up company to develop a chatbot meant to represent the artist Vincent van Gogh by ingesting his writings (Small, 2023). Beyond public figures, Project December (projectdecember.net/) is a venture that allows end-users to upload content from a deceased loved one (Fagone, 2021) in order to “simulate the dead” via a customized chatbot. The Augmented Eternity project at MIT is an academic endeavor that aims to allow people to create digital representations of themselves with the express purpose of being able to agentically represent them after death to members of their social network (Rahnama et al., 2021).
@@
While the examples shared above demonstrate the diversity of emerging approaches, a key trend in post-mortem AI is around griefbots. Griefbots (Jiménez-Alonso and de Luna, 2022; van der Vorst and Kamp, 2022), sometimes also called deathbots (Lindemann, 2022) or deadbots (Hollanek and Nowaczyk-Basińska, 2024), are typically described as chatbots (predating modern large language model technologies) designed to allow the bereaved to “converse” with the deceased, typically created posthumously by a third party who leverages general-purpose text the deceased created during their life. The aforementioned “Fredbot” and “Roman” would be considered examples of griefbots.
Because of their third-party origins, griefbots raise controversial ethical issues regarding privacy and consent (Grandinetti et al., 2020) which must be considered from multiple perspectives. For example, Hollanek and Nowaczyk‑Basińska (Hollanek and Nowaczyk-Basińska, 2024) emphasize the importance of considering the complex relationships among three classes of stakeholders for such bots: data donors, data recipients, and service interactants. These align with our own thinking, both in our design framework (see first- vs. third-party generative ghosts) and in our discussion of risks and benefits. However, we use the terms deceased and bereaved to emphasize interpersonal and relational aspects of clones, while also considering the impacts on society at large.
@@
Generative ghosts can be considered an extension of the concept of a griefbot, but not all griefbots would be considered generative ghosts. Generative ghosts go beyond the traditional definition of griefbots in several ways: they always include the ability to generate novel content in-character (whereas many griefbots only regurgitate content from a training corpus), potentially evolve over time, and possess agentic (Park et al., 2023; Shavit et al., 2023) capabilities such as the ability to participate in the economy or perform other complex tasks with limited oversight. Notably, our framework for generative ghosts envisions the possibility of first-party (rather than only third-party) creation as part of an end-of-life planning process. Also, generative ghosts may be designed to represent non-human entities (such as pets), and in many cases might exist pre-mortem as a generative clone. Finally, in contrast to griefbots, generative ghosts may be created for reasons other than to support the bereaved in their grieving process. The following section introduces more details about the design space of generative ghosts.
3. Generative Ghosts: A Design Space